A 100.0 kg football player moving at 8.0 m/s [N] smashes into a training dummy for 0.25 s and bounces back with a velocity of 6.0 m/s [S]. Find the average net force and average acceleration on the football player. 


1
Expert's answer
2021-02-22T10:25:41-0500
"F_{avg}\Delta t=m(v_f-v_i),""F_{avg}=\dfrac{m|(v_f-v_i)|}{\Delta t},""F_{avg}=\dfrac{100\ kg\cdot|(-6\ \dfrac{m}{s}-8\ \dfrac{m}{s})|}{0.25\ s}=5600\ N,""a_{avg}=\dfrac{F_{avg}}{m}=\dfrac{5600\ N}{100\ kg}=56\ \dfrac{m}{s^2}."